please help me solve this question :
find the values of 'a' and 'b' :
3+sqrt(2)/ 3-sqrt(2) = a+b sqrt(2)

Answer by rothauserc(4718)

3+sqrt(2)/ 3-sqrt(2) = a+b sqrt(2)
(3+sqrt(2)/ 3-sqrt(2)) * ((3+sqrt(2) / 3+sqrt(2)) =
(9+6sqrt(2)+2) / (9 - 2) =
(11+6sqrt(2) / 7) =
(11/7) + (6/7)sqrt(2), therefore
a = 11/7 and b = 6/7